# Step-by-Step Mastery: Enhancing Soft Constraint Following Ability of Large Language Models

### v2

@@ Abstract
It is crucial for large language models (LLMs) to follow instructions that involve multiple constraints. However, soft constraints are semantically related and difficult to verify through automated methods. These constraints remain a significant challenge for LLMs. To enhance the ability of LLMs to follow soft constraints, we initially design a pipeline to obtain high-quality outputs automatically. Additionally, to fully utilize the acquired data, we introduce a training paradigm based on curriculum learning. We experimentally evaluate the effectiveness of our methods in improving LLMs’ soft constraint following ability and analyze the factors driving the improvements. The datasets and code are publicly available at https://github.com/Rainier-rq/FollowSoftConstraints.
@@ 1 Introduction
In the application of LLMs, their instruction following ability is of paramount importance, especially when the instructions involve multiple constraints. Lou et al. (2024); Zeng et al. (2023); Zhou et al. (2023a). The capability of LLMs plays a critical role in aligning with human preferences, ensuring the reliability and helpfulness of the models’ outputs Wang et al. (2023b); Song et al. (2024).
@@
Following instructions with soft constraints is imperative for LLMs Jiang et al. (2023b); Qin et al. (2024). Existing work on improving the ability of LLMs to follow multiple constraints mainly focuses on hard constraints, which are typically based on structured data or fixed-format requirements He et al. (2024a). These constraints can be explicitly expressed as specific rules and directly verified through programming methods Zhou et al. (2023a). For example, Python can parse JSON to verify hard constraints. However, hard constraints fail to adequately capture the complexity in real-world scenarios shown in Fig. 1. Instructions in real-world applications often contain semantic-level limitations, which can be categorized as soft constraints. Soft constraints include restrictions related to content Liang et al. (2024); Zhang et al. (2023), specific backgrounds Shanahan et al. (2023); Liu et al. (2023), and stylistic objectives Sigurgeirsson and King (2024); Mukherjee et al. (2024). They are difficult to verify automatically through programming methods. A variety of tasks involve soft constraints, such as open-ended question answering Zhuang et al. (2023), role-playing Shanahan et al. (2023), and suggestion generation Baek et al. (2024). As shown in Fig. 1, following soft constraints is challenging for LLMs.
@@
Figure 1: In real-world scenarios, user instructions contain many soft constraints, posing challenges for LLMs. H and S denote hard constraints and soft constraints, respectively.
@@
Figure 2: The framework of our study. We first design a pipeline that automates the construction of datasets with high-quality outputs for soft constraint following. Then, we propose a method to utilize positive and negative samples generated during the pipeline. Finally, we introduce a new training paradigm that leverages curriculum learning to enhance LLMs’ soft constraint following ability. CL denotes curriculum Learning.
However, following soft constraints is a non-trivial task. First, existing research on soft constraints in LLMs mainly focuses on evaluation Chen et al. (2024a); Qin et al. (2024) rather than improving their following. Also, as shown in Fig. 1, soft constraints are ambiguous and challenging for LLMs in real applications Wang et al. (2024). They depend on subjective interpretations and specific contexts. Unlike hard constraints, they cannot be assessed with fixed rules or scripts. Soft constraint evaluation often relies on prompting LLMs, which involves various biases Wang et al. (2023a). The inherent difficulty makes it more challenging for LLMs to generalize from hard to soft constraints He et al. (2024a). Moreover, many studies utilize advanced models, such as GPT-4, to generate responses Xu et al. (2023); Chiang et al. (2023). Soft constraints also present significant challenges for these advanced models. On the FollowBench benchmark Jiang et al. (2023b), GPT-4 demonstrates a hard satisfaction rate of merely 74.4%, making the assurance of high-quality outputs difficult. However, many studies show that the quality of training data is more important than its quantity Zhou et al. (2024); Li et al. (2023a). Therefore, designing a more effective data construction pipeline is crucial.
@@
In this work, we systematically investigate strategies to enhance the ability of LLMs to follow instructions with soft constraints, with the framework shown in Fig. 2. When more constraints are added to the instruction, LLM struggles to learn to follow each constraint. Moreover, LLMs’ outputs may not follow all constraints. This inconsistency can hurt the quality of outputs. To address this, we progressively add constraints to construct multi-constraint instructions and incorporate Judger to reorder the outputs based on the extent of following the constraints to obtain high-quality outputs. To fully utilize both positive and negative outputs during the reordering process, we leverage the preference learning algorithm Direct Preference Optimization (DPO) Rafailov et al. (2024) as the training method. Subsequently, we propose a novel training paradigm that constructs a curriculum based on the number of constraints in the instruction. In this framework, the model increasingly learns how to make preference judgments, beginning with easier cases and moving towards more challenging ones. Our methods improve the model’s soft constraint following ability while maintaining general capabilities.
@@
Our contributions are summarized as follows: (1) We design a pipeline that automates the construction of datasets with high-quality outputs for soft constraint following. We also propose a method that utilizes positive and negative samples generated during the pipeline. (2) We introduce a new training paradigm that leverages curriculum learning to enhance LLMs’ soft constraint following ability. (3) We conduct extensive experiments to validate the effectiveness of our methods and analyze the reasons for the performance improvement.
@@ 2 Related Work
Soft Constraint Following Existing research on soft constraint following largely concentrates on evaluating the ability of LLMs to follow these constraints by constructing benchmarks Jiang et al. (2023b); Qin et al. (2024). These benchmarks typically include a variety of fine-grained constraint types Zhang et al. (2024), and the results from testing LLMs on these benchmarks suggest that LLMs often struggle to follow these constraints He et al. (2024b). Despite this, there is a notable paucity of research aimed at improving LLMs’ capacity to comply with soft constraints, especially for soft constraints. Soft constraints can be categorized into several types: (1) Content soft constraints involve restrictions on the scope or depth of the responses Zhou et al. (2023b); Ashok and Poczos (2024). (2) Situation soft constraints refer to the background limitations of the responses Wang et al. (2023c); Shao et al. (2023). (3) Style soft constraints limit the manner or tone of expressions Tao et al. (2024); Pu et al. (2024). Some works directly utilize responses generated by GPT-4 to construct datasets Sun et al. (2024); Peng et al. (2023). However, the responses to instructions with soft constraints are often unreliable. Different from these, our study focuses on how to construct datasets with high-quality outputs for improving LLMs’ soft constraint following ability.
Curriculum Learning Curriculum learning is a training strategy that mimics the learning process of humans by advancing from simpler to more complex tasks Soviany et al. (2022); Wang et al. (2021). Current research on LLMs’ curriculum learning can be broadly categorized into two primary paradigms: (1) Learning Based on Data Difficulty: This approach involves constructing curricula by ranking data according to various evaluation metrics. Metrics such as sequence length Pouransari et al. (2024), perplexity Liu et al. (2024) have been employed to guide this process. LLMs can also construct curricula through advanced planning Ryu et al. (2024). (2) Learning Based on Task Difficulty: This paradigm focuses on modifying the training tasks Chen et al. (2024b) or adjusting the training objectives Zhao et al. (2024b); Lee et al. (2024). However, our work organizes the curriculum based on the number of constraints in the instructions.
@@ 3 Method
In this section, we provide a detailed explanation of how to obtain high-quality outputs and how to leverage this data by establishing a new training paradigm. The pipeline is shown in Fig. 2.
3.1 High-quality Dataset
To enhance the ability of LLMs to follow soft constraints, we first construct a multi-constraint instruction following dataset. Existing works in dataset construction rely on advanced models to directly generate the outputs Sun et al. (2024). However, even GPT-4 is struggling to follow the instructions with complex constraints, especially when the instructions contain soft constraints which are more intractable Jiang et al. (2023b); Qin et al. (2024). To address this challenge, we design a pipeline to construct datasets with high-quality outputs for soft constraint following. This pipeline consists of two steps: progressive construction and Judger reordering.
@@ 3.1.1 Progressive Construction
To enable the model to learn how to follow each constraint, we propose a progressive dataset construction method. Specifically, we increase only one constraint at a time, enabling the model progressively learn to follow each constraint during the training process.
@@
Subsequently, we construct different types of soft constraints. Initially, we categorize the soft constraints into three types: content, situation, and style. Next, we randomly select 5 constraints for each seed instruction. For the soft constraints, LLMs are employed to generate corresponding descriptions. While for the hard constraints, descriptions are selected from a predefined list. The prompt used to construct soft constraints is detailed in the Appx. A.1.
To obtain multi-constraint instructions, we adopt a progressive construction approach. This approach is different from previous methods, which typically add all constraints at once, often making it challenging for the model to learn how to follow each constraint independently He et al. (2024a). We add only one constraint to the instruction at a time, allowing the model to focus on learning and mastering each constraint independently. This step-by-step process helps the model gradually adapt to the increasing complexity of the task, ensuring that it can effectively handle each constraint as it is introduced.
Specifically, for seed instruction $I_{0}$, we progressively add one constraint each time to form the instruction set $I=\{I_{1},I_{2},\dots,I_{n}\}$, where $n$ denotes the maximum number of constraints. For each instruction $I_{k}$ with $k$ constraints $(k=1,2,\dots,n)$, we use GPT-4o to generate the corresponding output $O_{k}=\text{LLM}(I_{k})$. After performing inference on all the instructions in the instruction set $I$, we obtain the output set $O=\{O_{1},O_{2},\dots,O_{n}\}$.
@@ 3.1.2 Judger Reordering
In §3.1.1, we progressively increase the constraints, but the quality of the outputs may not improve incrementally. To address this, we introduce Judger to reorder the outputs based on the extent of constraint following to ensure the quality of outputs.
During the progressive construction process in §3.1.1, as new constraints are continuously added, the model’s responses may overlook previously incorporated constraints, leading to a decrease in constraint following. To obtain high-quality outputs, we introduce Judger, where LLM is prompted to compare two outputs before and after adding the new constraint, to determine which better follows the updated instruction. The two outputs in each comparison are recorded, and the one deemed better by Judger is used for the next round of comparison. By iteratively ranking the outputs, the constructed data is consistent with constraint following, thereby improving the output quality.
@@
In each ranking, we can obtain the output $O_{w_{k}}$ which follows the current instruction $I_{k}$ better and the output $O_{l_{k}}$ which follows less. Finally, after completing all $n$ rankings, we obtain the positive set $O_{w}=\{O_{w_{1}},O_{w_{2}},\dots,O_{w_{n}}\}$, which consists of outputs that follow their respective instructions better. We also obtain the negative set $O_{l}=\{O_{l_{1}},O_{l_{2}},\dots,O_{l_{n}}\}$, which contains outputs that less follow. The prompt used to reorder outputs and cases are detailed in the Appx. A.2.
3.2 Curriculum-based Training Paradigm
In §3.1.2, we use Judger to obtain the positive set $O_{w}$ and the negative set $O_{l}$. Supervised Fine-Tuning (SFT) Ouyang et al. (2022) only uses the positive samples to train the model. However, the negative samples also contain valuable supervision information. Hence, we adopt reinforcement learning Rafailov et al. (2024) to leverage both the positive and negative sets. Moreover, we develop a training paradigm based on curriculum learning to enhance the training process.
Given the positive set and the negative set, we can construct the training dataset with $n$ triplets: ($I_{1}$, $O_{w_{1}}$, $O_{l_{1}}$), ($I_{2}$, $O_{w_{2}}$, $O_{l_{2}}$), …, ($I_{n}$, $O_{w_{n}}$, $O_{l_{n}}$). In each triplet, the output from $O_{w}$ is preferred than the output from $O_{l}$. To model this preference relationship, we apply Direct Preference Optimization (DPO) Rafailov et al. (2024) as the training method.
Additionally, in the DPO training process, the model is required to learn preference judgments. As the number of constraints in the instruction increases, the complexity of judgments also rises. Inspired by curriculum learning Bengio et al. (2009), we propose a curriculum learning training approach for preference learning, starting with simpler preference judgments involving fewer constraints and progressively advancing to more complex judgments involving more constraints.
[…]
Based on the preference data and the curriculum-based training paradigm, the loss function of DPO training can be defined as follows:
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-{\mathbb{E}}_{(I_{k},O_{w_{k}},O_{l_{k}})\sim\mathcal{D}}[\text{log}\sigma(\beta\text{log}\frac{\pi_{\theta}(O_{w_{k}}|I_{k})}{\pi_{\text{ref}}(O_{w_{k}}|I_{k})}$
@@ 3.2 Curriculum-based Training Paradigm
where $pi_{\theta}$ represents the current model, and $pi_{\text{ref}}$ denotes the reference model.
@@
$\mathcal{L}_{\text{SFT}}(\pi_{\theta})=-\mathbb{E}_{(I_{k},O_{w_{k}})\sim D}[\log\pi_{\theta}(O_{w_{k}}|I_{k})].$
3.3 Dataset Statistics
3.3.1 Diversity
To show the diversity of our dataset, we analyze the verb-noun structure of data. As shown in Fig. 3, we illustrate the top 10 verbs in the inner circle and their 3 most frequent direct noun objects in the outer circle. The result reveals the instructions in our dataset encompass a diverse set of linguistic patterns. This diversity of our dataset is crucial for enhancing the model’s ability to generalize across various types of constraints.
Method Data Quality
Nums. Cons. Pair. Open.
Conifer Sun et al. (2024) 13600 H/S $\times$ $\checkmark$
Suri Pham et al. (2024) 10000 S $\times$ $\checkmark$

### v3

@@ Abstract
It is crucial for large language models (LLMs) to follow instructions that involve multiple constraints. However, it is an unexplored area to enhance LLMs’ ability to follow soft constraints. To bridge the gap, we initially design a pipeline to construct datasets with high-quality outputs automatically. Additionally, to fully utilize the positive and negative samples generated during the data construction process, we choose Direct Preference Optimization (DPO) as the training method. Furthermore, taking into account the difficulty of soft constraints indicated by the number of constraints, we design a curriculum learning training paradigm based on the constraint quantity. We experimentally evaluate the effectiveness of our methods in improving LLMs’ soft constraint following ability and analyze the factors driving the improvements. The datasets and code are publicly available at https://github.com/Rainier-rq/FollowSoftConstraint.
@@ 1 Introduction
In the application of LLMs, the instruction following ability is of paramount importance, especially when the instructions involve multiple constraints Lou et al. (2024); Zeng et al. (2023); Zhou et al. (2023a). The capability of LLMs plays a critical role in aligning with human preferences, ensuring the reliability and helpfulness of the models’ outputs Wang et al. (2023a); Song et al. (2024).
@@
Following instructions with soft constraints is imperative for LLMs Jiang et al. (2023b); Qin et al. (2024). Constraints can be categorized into soft and hard constraints. Hard constraints can be explicitly expressed as specific rules and directly verified through programming methods Zhou et al. (2023a); He et al. (2024a). For example, Python can parse JSON data to verify whether it follows specific format constraints. However, instructions in real-world applications often contain semantic-level limitations, which can be categorized as soft constraints. Soft constraints include restrictions related to content Liang et al. (2024); Zhang et al. (2023), specific backgrounds Shanahan et al. (2023); Liu et al. (2023), and the style of expression Sigurgeirsson and King (2024); Mukherjee et al. (2024). They are difficult to verify automatically through programming methods. As shown in Fig. 1, following soft constraints is challenging for LLMs.
@@
Figure 1: In real-world scenarios, user instructions contain many soft constraints, posing challenges for LLMs. We use bold to represent soft constraints.
@@
Figure 2: The framework of our study. We first design a pipeline to automatically construct datasets with high-quality outputs for soft constraint following. Then, we propose a DPO training paradigm based on curriculum learning. CL denotes curriculum learning.
Improving LLMs’ soft constraint following ability is under-explored. Much of the existing work focuses on hard constraints, and it is challenging for LLMs to generalize from hard to soft constraints He et al. (2024a). Additionally, existing research focuses on evaluating LLMs’ soft constraint following ability Chen et al. (2024a); Qin et al. (2024) rather than improving the ability. Overall, existing work on soft constraints has following three limitations: First, directly using advanced models to generate outputs for constructing the dataset results in low quality Xu et al. (2023); Chiang et al. (2023). Soft constraints are especially challenging for these models. On FollowBench Jiang et al. (2023b), GPT-4 demonstrates a hard satisfaction rate of merely 74.4%, making the assurance of high-quality outputs difficult. Second, simply using Supervised Fine-Tuning (SFT) Ouyang et al. (2022) to train the model leads to poor efficiency and performance. Slight changes about the soft constraint in the instruction can result in significant variations in the output He et al. (2024a). It is hard to learn these differences just from positive samples. Third, existing work ignores the difficulty of soft constraints indicated by the number of constraints. Many studies show that organizing training data in the order of the curriculum is better than random shuffling Sun et al. (2024); Lee et al. (2024). Therefore, an efficient training paradigm is required.
@@
In this work, we systematically investigate strategies to enhance the ability of LLMs to follow soft constraints, with the framework shown in Fig. 2, including constructing datasets with high-quality outputs and proposing a DPO training paradigm based on curriculum learning. To construct datasets with high-quality outputs, we progressively add constraints to the instructions and incorporate Judger to reorder the outputs based on constraint following. To fully leverage the positive and negative samples generated during Judger reordering, we apply Direct Preference Optimization (DPO) Rafailov et al. (2024) as the training method. By contrasting positive and negative outputs, it can capture how minor soft constraint adjustments affect output preferences. To account for the difficulty of soft constraints indicated by the number of constraints, we propose a training paradigm that constructs curricula based on the number of constraints in the instruction. In this paradigm, the model starts by learning simpler tasks (fewer constraints) and progresses to more complex ones (more constraints). Our methods can improve LLMs’ soft constraint following ability effectively.
@@
Our contributions are summarized as follows: (1) We design a pipeline to automatically construct datasets with high-quality outputs for soft constraint following. (2) We introduce a DPO training paradigm that constructs curricula based on the number of constraints. (3) We conduct extensive experiments to validate the effectiveness of our methods and analyze the reasons for the performance improvements.
@@ 2 Related Work
Soft Constraint Following Existing research on soft constraint following focuses on evaluating the ability of LLMs by constructing benchmarks Jiang et al. (2023b); Qin et al. (2024). These benchmarks include a variety of fine-grained constraint types Zhang et al. (2024). These constraints can be categorized into several types: (1) Content soft constraints involve restrictions on the scope or depth of the responses Zhou et al. (2023b); Ashok and Poczos (2024). (2) Situation soft constraints refer to the background limitations Wang et al. (2023b); Shao et al. (2023). (3) Style soft constraints limit the manner of expressions Tao et al. (2024); Pu et al. (2024). Some works directly utilize responses generated by GPT-4 to construct datasets Sun et al. (2024); Peng et al. (2023). However, LLMs struggle to follow soft constraints He et al. (2024b), making responses unreliable. Different from these, we construct datasets with high-quality outputs to improve LLMs’ soft constraint following ability.
Curriculum Learning Curriculum learning is a training strategy that mimics the learning process of humans from simpler to more complex tasks Soviany et al. (2022); Wang et al. (2021). Current research on LLMs’ curriculum learning can be categorized into two primary paradigms: (1) Learning Based on Data Difficulty: This approach organizes the training data sequence according to various evaluation metrics. Metrics such as sequence length Pouransari et al. (2024), perplexity Liu et al. (2024) have been employed to guide this process. LLMs can also construct curricula by organizing the training data sequence in a strategic way Ryu et al. (2024). (2) Learning Based on Task Difficulty: This paradigm focuses on changing the training tasks Chen et al. (2024b) or adjusting the training objectives Zhao et al. (2024b); Lee et al. (2024). However, our work organizes the curriculum based on the number of constraints .
@@ 3 Method
In this section, we provide a detailed explanation of how to obtain datasets with high-quality outputs and how to leverage the dataset by establishing a curriculum learning training paradigm. The framework is shown in Fig. 2.
3.1 Dataset Construction
We first construct a multi-constraint instruction following dataset. Adding all constraints at once increases the complexity of the instructions rapidly, making it difficult for the model to understand all constraints He et al. (2024a). To addresss this, we propose a progressive construction method, adding one constraint at a time, which allows the model to understand and learn to follow each constraint effectively. Moreover, existing works in dataset construction rely on advanced models to directly generate the outputs Sun et al. (2024). However, even GPT-4 is struggling to follow the instructions with complex soft constraints Jiang et al. (2023b); Qin et al. (2024). To obtain high-quality outputs, we use Judger to reorder the outputs based on the extent of constraint following. Overall, our pipeline consists of two successive steps: Progressive Construction and Judger Reordering.
@@ 3.1.1 Progressive Construction
To enable the model to effectively learn how to follow each constraint, we propose a progressive construction method. Specifically, we add only one constraint at a time, enabling the model progressively learn to follow each constraint during the training process.
@@
Subsequently, we construct different types of soft constraints. Initially, we categorize the soft constraints into three types: content, situation, and style. Next, we randomly select constraints for each seed instruction. For the soft constraint type we select, GPT-4o is employed to generate corresponding descriptions. The prompt used to construct soft constraints is detailed in the Appx. A.1. For the hard constraint type, we select the description from a predefined description list.
To obtain multi-constraint instructions, we adopt a progressive construction approach. As shown in Fig. 2, we add only one constraint to the instruction at a time, allowing the model to focus on learning to follow each constraint. This process helps the model gradually adapt to the increasing complexity of the constraint following task and balance multiple constraints effectively. Specifically, for seed instruction $I_{0}$, we progressively add one constraint each time to form the instruction set $I=\{I_{1},I_{2},\dots,I_{n}\}$, where $n$ denotes the maximum number of constraints. For each instruction $I_{k}$ with $k$ constraints $(k=1,2,\dots,n)$, we use GPT-4o to generate the corresponding output $O_{k}=\text{LLM}(I_{k})$. After performing inference on all the instructions in the instruction set $I$, we obtain the output set $O=\{O_{1},O_{2},\dots,O_{n}\}$.
@@ 3.1.2 Judger Reordering
In §3.1.1, we progressively increase the constraints, but the quality of the outputs may not improve incrementally. To address this, we introduce Judger to reorder the outputs based on constraint following to ensure the quality of outputs.
During the progressive construction process in §3.1.1, as new constraints are progressively added, the model’s responses may overlook previously added constraints, leading to a decrease in the output quality. As shown in Fig. 2, to obtain high-quality outputs, we introduce Judger, where GPT-4o is prompted to compare two outputs before and after adding the new constraint, to determine which better follows the updated instruction. The two outputs in each comparison are recorded, and the one deemed better by Judger is used for the next round of comparison. By iteratively ranking the outputs, the constructed data is consistent with constraint following, thereby improving the output quality.
@@
In each ranking, we can obtain the output $O_{w_{k}}$ which follows the current instruction $I_{k}$ better and the output $O_{l_{k}}$ which follows less. Finally, after completing all $n$ rankings, we obtain the positive set $O_{w}=\{O_{w_{1}},O_{w_{2}},\dots,O_{w_{n}}\}$, which consists of outputs that follow their respective instructions better. We also obtain the negative set $O_{l}=\{O_{l_{1}},O_{l_{2}},\dots,O_{l_{n}}\}$, which contains outputs that less follow. The prompt used to reorder outputs and reordering cases are detailed in the Appx. A.2.
3.2 Curriculum Learning Training Paradigm
In §3.1.2, we use Judger to obtain the positive set $O_{w}$ and the negative set $O_{l}$. As shown in Fig. 2, to fully leverage both the positive and negative sets, we apply DPO as the training method. To account for the difficulty of soft constraints indicated by the number of constraints , we establish a curriculum learning training paradigm based on the number of constraints in the instruction.
Given the positive set and the negative set, we can construct the training dataset with $n$ triplets: ($I_{1}$, $O_{w_{1}}$, $O_{l_{1}}$), ($I_{2}$, $O_{w_{2}}$, $O_{l_{2}}$), …, ($I_{n}$, $O_{w_{n}}$, $O_{l_{n}}$). In each triplet, the output from $O_{w}$ is preferred than the output from $O_{l}$. To fully leverage these samples, we apply DPO to train the model. To enable the model to learn from easy to hard tasks during training, we propose a curriculum learning training paradigm based on the number of constraints, starting with simpler tasks involving fewer soft constraints and progressively advancing to more complex tasks involving more soft constraints.
Specifically, for curriculum $k$, the training dataset $D_{k}$ contains the triplet $(I_{k},O_{w_{k}},O_{l_{k}})$, where $I_{k}$ represents the instruction with $k$ constraints, and $O_{w_{k}}$ and $O_{l_{k}}$ denote the corresponding outputs in preference learning. The model begins with the simplest dataset, $D_{1}$, and sequentially progresses through $D_{2}$ to $D_{n}$, gradually enhancing its ability to handle more soft constraints. The complete curriculum is defined as $D=\{D_{1},D_{2},\dots,D_{n}\}$. This stepwise approach ensures that the model first builds a strong foundation by learning from simpler datasets with fewer soft constraints, and gradually adapts to more complex datasets with more constraints. To prevent catastrophic forgetting during training McCloskey and Cohen (1989), we mix each curriculum’s data with a proportion of 10k ShareGPT examples based on its data size Chiang et al. (2023). Based on the curriculum learning training paradigm, the loss function of DPO training is as follows:
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-{\mathbb{E}}_{(I_{k},O_{w_{k}},O_{l_{k}})\sim{D}_{k}}[\text{log}\sigma(\beta\text{log}\frac{\pi_{\theta}(O_{w_{k}}|I_{k})}{\pi_{\text{ref}}(O_{w_{k}}|I_{k})}$
@@ 3.2 Curriculum Learning Training Paradigm
where $\pi_{\theta}$ represents the current model, and $\pi_{\text{ref}}$ denotes the reference model.
@@
$\mathcal{L}_{\text{SFT}}(\pi_{\theta})=-\mathbb{E}_{(I_{k},O_{w_{k}})\sim D_{k}}[\log\pi_{\theta}(O_{w_{k}}|I_{k})].$
3.3 Analysis and Comparison
Curriculum # Constraints # Preference Pairs Avg Length
Curri.1 1 3714 369
Curri.2 2 3494 422
Curri.3 3 3387 461
Curri.4 4 3300 503
Curri.5 5 3148 516
Table 1: Statistics of curricula. # Constraints refers to the number of constraints in each instruction. # Preference Pairs refers to the number of preference pairs. “Avg Length” denotes the average instruction length.
Figure 3: The verb frequency in the instructions.
3.3.1 Data Statistics
We present a statistical analysis of different curricula in Tab. 1. The results show that the number of constraints and instruction length in the curriculum continuously increase. Each curriculum contains a large scale of preference data. To show the diversity of our dataset, we analyze the frequency of verbs in the instructions. As shown in Fig. 3, the instructions contain a variety of verbs, reflecting diverse linguistic patterns. This diversity is crucial for enhancing the model’s ability to generalize across different types of constraints.
Method Nums. Cons. Pair. Prog.
Conifer Sun et al. (2024) 13606 H/S $\times$ $\checkmark$
Suri Pham et al. (2024) 20000 S $\checkmark$ $\times$